OpenStudy (love10129151):

Task 2 Write an equation in slope-intercept, point-slope, or standard form for the line with the given information. Explain why you chose the form you used. a. passes through (-1, 4) and (-5, 2) Equation: _____________________________ Explain why you chose the form you used. ________________________________________ ________________________________________________________ b. slope 2, y-intercept -4 Equation: _____________________________ Explain why you chose the form you used.

OpenStudy (love10129151):

i need help now

zepdrix (zepdrix):

hi c:

OpenStudy (love10129151):

hi

OpenStudy (love10129151):

can you help me

zepdrix (zepdrix):

For the first one, since we're given a coordinate pair, let's use the `point-slope` form of a line.\[\Large y-y_1=m(x-x_1)\]Where we'll use the coordinate pair, \(\Large (-1,\;4)\) as our \(\Large (x_1,\;y_1)\).

zepdrix (zepdrix):

So all we need to do is find the slope, \(\Large m\), of the line.

zepdrix (zepdrix):

\[\Large m\quad=\quad \frac{y_2-y_1}{x_2-x_1}\] And our coordinate pairs are,\[\Large (x_1,\;y_1)\quad=\quad (-1,\;4)\]\[\Large (x_2,\;y_2)\quad=\quad (-5,2)\]

zepdrix (zepdrix):

Understand how to find the slope using the coordinate pairs? :o

OpenStudy (love10129151):

yes

OpenStudy (love10129151):

can you help me with the second one

zepdrix (zepdrix):

For the second one, since were given the `slope` and the `y-intercept` it makes sense that we should use the `slope-intercept` form of a line.\[\Large y\quad=\quad mx+b\]Where the slope is \(\Large m\) and the y-intercept is \(\Large b\).

zepdrix (zepdrix):

\[\Large\bf y\quad=\quad \color{#008353 }{m}x+\color{#DD4747}{b}\]We are told that the slope is \(\Large\bf \color{#008353 }{2}\) and the y-intercept is \(\Large\bf \color{#DD4747 }{-4}\). Understand where to plug those values in? :) Hopefully the color helps :x lol

OpenStudy (love10129151):

ok can you help me with two more

OpenStudy (love10129151):

c. has an x-intercept of 6 and a y-intercept of 3 Equation: _____________________________ Explain why you chose the form you used. ________________________________________ ________________________________________________________ d. passes through (1, 2) with slope -5/3 . Equation: _____________________________ Explain why you chose the form you used. ________________________________________ ________________________________________________________

zepdrix (zepdrix):

Hmm for the first one, I'm not very familiar with using `Standard Form` but I assume that's what they want us to do. x-intercept of 6 tells us that x=6 when y is zero.\[\Large x+0=6\]And a y-intercept of 3 tells us that y=3 when x is zero.\[\Large 0+y=3\qquad\to\qquad 0+2y=6\] Which gives us uhhh:\[\Large x+2y=6\]Mmm I dunno I can't think of a good way to explain that one :\ Standard Form is weird :p

zepdrix (zepdrix):

For the last problem, we were given a `point` and a `slope`. So we should use the `point-slope` form of a line.\[\Large y-y_1\quad=\quad m(x-x_1)\]Where our given point is:\[\Large (x_1,\;y_1)\quad=\quad (1,\;2)\]and slope,\[\Large m\quad=\quad-\frac{5}{3}\]Understand how to plug them into the formula?

OpenStudy (love10129151):

yes can you help me with a graph

OpenStudy (love10129151):



OpenStudy (love10129151):

Rate of change: ___________ How did you find the rate of change? ___________________________________